# Relative Energy Learning for LiDAR Out-of-Distribution Detection

### v1

@@ 3.3 Relative Energy Learning
REL introduces a contrastive energy margin between ID and negative OOD logits, yielding a shift-invariant energy gap between ID and OOD samples, and employs an imbalance-aware loss with a single weighting parameter. This formulation eliminates the need for delicate hyperparameter tuning while explicitly leveraging scarce auxiliary OOD samples to sharpen decision boundaries under class imbalance.
@@ 3.4 Point Raise: A Lightweight Auxiliary OOD Synthesis Algorithm
As shown in Algorithm 1, the proposed Point Raise algorithm first samples a random road point, gathers its neighbors within a random radius using a KD-tree, contracts them radially toward the sensor with an adaptive decay scaled by pull factor $\gamma$, and finally perturbs their heights by random offsets in $h_{\min},h_{\max}]$. The modified points are relabeled as a reserved OOD class.
@@
Each point is then scaled by $s=\exp(-a\cdot d_{\text{shift}})$, so that near points remain almost unchanged while far points are pulled inward, compacting the cluster into an object-like anomaly.
@@
AUROC $\uparrow$ FPR@95 $\downarrow$ AP $\uparrow$ RecallQ SQ RQ UQ PQ
@@
AUROC $\uparrow$ FPR@95 $\downarrow$ AP $\uparrow$ RecallQ SQ RQ UQ PQ
@@ 4.3 Implementation Details
We initialize the model from a Mask4Former checkpoint pretrained on SemanticKITTI [behley2019semantickitti] and Panoptic CUDAL [tseng2025panopticcudal], and apply the proposed Point Raise and Relative Energy Learning (REL) techniques using only a single additional sequence from the STU closed-set training set.
@@ Acknowledgments
The first two authors acknowledge the financial support from The University of Melbourne through the Melbourne Research Scholarship. This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.
@@ 7 Additional Qualitative Results
Table 8: Closed-set per-class performance (Panoptic Quality) of the methods on validation sets of STU [nekrasov2025stu] dataset. With REL, our model retains comparable closed-set panoptic segmentation performance to standard Mask4Former-3D training, while largely outperforming OOD-specific methods such as the void classifier on OOD detection.

### v2

@@ 3.3 Relative Energy Learning
REL introduces a contrastive energy margin between ID and negative OOD logits, yielding a shift-invariant energy gap between ID and OOD samples, and employs an imbalance-aware loss with a single weighting parameter. Introducing negative logits helps to better partition the logit space and explicitly model OOD samples. This formulation eliminates the need for delicate hyperparameter tuning while explicitly leveraging scarce auxiliary OOD samples to sharpen decision boundaries under class imbalance.
@@ 3.4 Point Raise: A Lightweight Auxiliary OOD Synthesis Algorithm
As shown in Algorithm 1, the proposed Point Raise algorithm first samples a random road point, gathers its neighbors within a random radius using a KD-tree, contracts them radially toward the sensor with an adaptive decay scaled by the pull factor $\gamma$, and finally perturbs their heights by random offsets in $h_{\min},h_{\max}]$. The modified points are relabeled as a reserved OOD class.
@@
Each point is then scaled by $s=\exp(-a\cdot d_{\text{shift}})$, so that the near points remain almost unchanged while the far points are pulled inward, compacting the cluster into an object-like anomaly.
@@
AUROC $\uparrow$ FPR@95 $\downarrow$ AP $\uparrow$ RecallQ $\uparrow$ SQ $\uparrow$ RQ $\uparrow$ UQ $\uparrow$ PQ $\uparrow$
@@
AUROC $\uparrow$ FPR@95 $\downarrow$ AP $\uparrow$ RecallQ $\uparrow$ SQ $\uparrow$ RQ $\uparrow$ UQ $\uparrow$ PQ $\uparrow$
@@ 4.3 Implementation Details
We initialize the model from a Mask4Former checkpoint pretrained on SemanticKITTI [behley2019semantickitti] and Panoptic-CUDAL [tseng2025panopticcudal], and apply the proposed Point Raise and Relative Energy Learning (REL) techniques using only a single additional sequence from the STU closed-set training set.
@@ Acknowledgments
The first two authors acknowledge the financial support from The University of Melbourne through the Melbourne Research Scholarship. This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.
@@ 7 Additional Qualitative Results
Table 8: Closed-set per-class performance (Panoptic Quality) of the methods on validation sets of STU [nekrasov2025stu] dataset. With REL, our model retains comparable closed-set panoptic segmentation performance to standard Mask4Former-3D training, while largely outperforming OOD-specific methods such as the void classifier on OOD detection. The bicycle class is severely underrepresented in STU, and fine-tuning on STU leads to a noticeable drop in PQ.